1) Sasha owns two investments, A and B, that have a combined total value of 42,100 dollars. Investment A is expected to pay 27,400 dollars in 3 year(s) from today and has an expected return of 3.59 percent per year. Investment B is expected to pay 51,075 dollars in T years from today and has an expected return of 9.17 percent per year. What is T, the number of years from today that investment B is expected to pay 51,075 dollars? Round your answer to 2 decimal places (for example, 2.89, 14.70, or 6.00).

2) 1 year(s) ago, Mack invested 5,630 dollars. In 2 year(s) from today, he expects to have 8,730 dollars. If Mack expects to earn the same annual return after 2 year(s) from today as the annual rate implied from the past and expected values given in the problem, then how much does Mack expect to have in 7 years from today?

3) 1 year(s) ago, Isaac had 243,500 dollars in his account. In 6 year(s), he expects to have 376,400 dollars. If he has earned and expects to earn the same return each year from 1 year(s) ago to 6 year(s) from today, then how much does he expect to have in 3 year(s) from today?

4) 2 year(s) ago, Fatima invested 6,050 dollars. In 1 year(s) from today, she expects to have 7,560 dollars. If Fatima expects to earn the same annual return after 1 year(s) from today as the annual rate implied from the past and expected values given in the problem, then in how many years from today does she expect to have exactly 10,920 dollars? Round your answer to 2 decimal places (for example, 2.89, 14.70, or 6.00)
Answer #1

1)

combined value=42100

27400/(1+3.59%)^3+51075/(1+9.17%)^T=42100

51075/(1+9.17%)^T=(42100-(27400/(1+3.59%)^3))

1.0917^T=51075/(42100-(27400/(1+3.59%)^3))

1.0917^T=2.92675

take natural log on both sides

T=ln(2.92675)/ln(1.0917)

=12.24 years

the above is answer..

we do only one question based on Chegg rule